# DualToken: Towards Unifying Visual Understanding and Generation with Dual Visual Vocabularies

### v1

@@ Abstract
The differing representation spaces required for visual understanding and generation pose a challenge in unifying them within the autoregressive paradigm of large language models. A vision tokenizer trained for reconstruction excels at capturing low-level perceptual details, making it well-suited for visual generation but lacking high-level semantic representations for understanding tasks. Conversely, a vision encoder trained via contrastive learning aligns well with language but struggles to decode back into the pixel space for generation tasks. To bridge this gap, we propose DualToken, a method that unifies representations for both understanding and generation within a single tokenizer. However, directly integrating reconstruction and semantic objectives in a single tokenizer creates conflicts, leading to degraded performance in both reconstruction quality and semantic performance. Instead of forcing a single codebook to handle both semantic and perceptual information, DualToken disentangles them by introducing separate codebooks for high and low-level features, effectively transforming their inherent conflict into a synergistic relationship. As a result, DualToken achieves state-of-the-art performance in both reconstruction and semantic tasks while demonstrating remarkable effectiveness in downstream MLLM understanding and generation tasks. Notably, we also show that DualToken, as a unified tokenizer, surpasses the naive combination of two distinct types vision encoders, providing superior performance within a unified MLLM.
@@ 2 Related Work
Unified Multimodal Large Language Models To integrate visual understanding and generation within a single MLLM, a common approach is to externally connect an LLM with a Diffusion Model [38, 8, 12, 13]. While these models can perform both vision-language understanding and visual generation tasks, they essentially combine two separate models rather than forming a truly unified framework. In contrast, pure autoregressive (AR) models, such as Chameleon [53, 39] and Emu3 [44], have demonstrated the feasibility of jointly modeling vision and language through a unified next-token prediction objective. Specifically, visual inputs are first discretized into visual tokens with a vision tokenizer. These visual tokens are then interleaved with text tokens to construct a multimodal sequence. This approach provides an elegant, fully end-to-end solution that unifies visual understanding and generation within the same autoregressive paradigm. However, pure AR architectures like Chameleon introduce generative capabilities at the cost of significantly weaker visual understanding compared to models specifically designed for understanding tasks. An empirical explanation for this [46, 49] is that models like Chameleon often utilize VQVAE [42, 10] encoders as their vision tokenizers. Since VQVAE is trained with a reconstruction objective, its encoder primarily extracts low-level visual features optimized for generation rather than the high-level semantic representations required for vision-language understanding. Therefore, improving the vision-language understanding performance of AR-based models necessitates a vision tokenizer that effectively captures both low-level representations for generation and high-level semantic representations for understanding.
Recent research has actively explored solutions in this direction. VILA-U [46] and Muse-VL [49] attempt to integrate both reconstruction and semantic objectives, aiming to develop a unified tokenizer through joint training. However, due to the inherent disparity between semantic and texture features, these models struggle to strike an optimal balance between the two objectives, resulting in subpar performance in both visual understanding and generation. Similar to ours, TokenFlow [33] introduces separate codebooks with a shared-mapping mechanism. However, there are key differences: first, TokenFlow employs separate vision towers to extract semantic and low-level features rather than using a unified architecture. Second, the shared IDs obtained through the shared-mapping mechanism may not be the optimal matches for either semantics or texture, potentially introducing additional losses for both types of information. Moreover, TokenFlow has not yet proposed a fully unified model for both generation and understanding; instead, it trains separate models dedicated to these tasks.
@@ 3.1 Motivation
To validate this viewpoint, we started by a preliminary experiment following the LLaVA-1.5 pipeline [26]. As shown in Table 2, compared to the original SigLIP model, encoders trained with reconstruction objective exhibit a significant performance drop in downstream MLLM vision-language understanding tasks, validating that high-level semantic features are more critical for visual reasoning in MLLMs than low-level perceptual features.
However, to achieve both visual understanding and generation within a single MLLM, it is essential to decode the visual tokens back into pixel space as accurately as possible. However, since the SigLIP encoder focuses on high-level semantic information and lacks texture details, simply discretizing its features and training a decoder without tuning the encoder results in poor image reconstruction quality. Therefore, proposing a unified tokenizer is crucial to enable high-quality visual understanding and generation within a singe MLLM.
@@ 3.2 Unified Vision Tokenizer with Dual Codebooks
We utilize an Exponential Moving Average (EMA) strategy [35, 41] to update the codebook. To enhance codebook utilization, we implement a restart strategy where a cluster within the codebook is randomly replaced with an input from the current batch if it remains unused for a certain number of steps. To ensure the encoder outputs align closely with the codebook entries, we utilize a Vector Quantization (VQ) commitment loss. The VQ commitment loss is defined as:

### v2

@@ Abstract
The differing representation spaces required for visual understanding and generation pose a challenge in unifying them within the autoregressive paradigm of large language models. A vision tokenizer trained for reconstruction excels at capturing low-level perceptual details, making it well-suited for visual generation but lacking high-level semantic representations for understanding tasks. Conversely, a vision encoder trained via contrastive learning aligns well with language but struggles to decode back into the pixel space for generation tasks. To bridge this gap, we propose DualToken, a method that unifies representations for both understanding and generation within a single tokenizer. However, directly integrating reconstruction and semantic objectives in a single tokenizer creates conflicts, leading to degraded performance in both reconstruction quality and semantic performance. Instead of forcing a single codebook to handle both semantic and perceptual information, DualToken disentangles them by introducing separate codebooks for high and low-level features, effectively transforming their inherent conflict into a synergistic relationship. As a result, DualToken achieves state-of-the-art performance in both reconstruction and semantic tasks while demonstrating remarkable effectiveness in downstream MLLM understanding and generation tasks. Notably, we also show that DualToken, as a unified tokenizer, surpasses the naive combination of two distinct types vision encoders, providing superior performance within a unified MLLM. The code and models will be available at https://github.com/songweii/DualToken.
@@ 1 Introduction
FQGAN-Dual ($\downarrow$16) [3] 0.94 22.02 -
FQGAN-Triple ($\downarrow$16) [3] 0.76 22.73 -
@@ 2 Related Work
Unified Multimodal Large Language Models A common strategy for integrating visual understanding and generation within a single MLLM is to externally connect an LLM with a Diffusion Model [39, 9, 13, 14]. However, pure autoregressive (AR) architectures offer a more elegant, fully end-to-end solution by unifying both tasks within the same autoregressive framework. Representative models like Chameleon [54, 40] and Emu3 [45], have demonstrated the feasibility of jointly modeling vision and language through a unified next-token prediction objective. Specifically, visual inputs are first discretized into visual tokens with a vision tokenizer. These visual tokens are then interleaved with text tokens to construct a multimodal sequence. However, pure AR architectures like Chameleon introduce generative capabilities at the cost of significantly weaker visual understanding compared to models specifically designed for understanding tasks. An empirical explanation for this [47, 50] is that models like Chameleon typically utilize VQVAE [43, 11] encoders as their vision tokenizers. Since VQVAE is trained with a reconstruction objective, its encoder primarily extracts low-level visual features optimized for generation rather than the high-level semantic representations required for vision-language understanding. Therefore, improving the vision-language understanding performance of AR-based models necessitates a vision tokenizer that effectively captures both low-level representations for generation and high-level semantic representations for understanding.
Recent research has actively explored solutions in this direction. VILA-U [47] and MUSE-VL [50] strive to build a unified tokenizer by jointly training on both reconstruction and semantic objectives. However, due to the inherent disparity between semantic and texture features, they struggle to strike an optimal balance between the two objectives, resulting in subpar performance in both tasks. As discussed in FQGAN [3], decomposing the codebook in a divide-and-conquer manner may offer a more fundamental solution to this conflict. TokenFlow [34] employs separate codebooks with a shared-mapping mechanism. However, key differences set our approach apart. First, TokenFlow relies on distinct vision towers to extract semantic and low-level features, rather than leveraging a unified architecture. Second, the shared IDs obtained through the shared-mapping mechanism may not be the optimal matches for either semantics or texture, potentially introducing additional losses in both domains. TokenFlow has yet to demonstrate its effectiveness in a unified MLLM setting that supports both generation and understanding; instead, it trains separate models dedicated to each task.
@@ 3.1 Motivation
To validate this viewpoint, we started by a preliminary experiment following the LLaVA-1.5 pipeline [27]. As shown in Table 2, compared to the original SigLIP model, encoders trained with reconstruction objective exhibit a significant drop in downstream MLLM vision-language understanding performance, validating that high-level semantic features are more critical for visual reasoning in MLLMs than low-level perceptual features. However, to achieve both visual understanding and generation within a single MLLM, it is essential to decode the visual tokens back into pixel space as accurately as possible. However, since the SigLIP encoder focuses on high-level semantic information rather than texture details, simply discretizing its features and training a decoder without tuning the encoder results in poor image reconstruction quality. Therefore, proposing a unified tokenizer is crucial to enable high-quality visual understanding and generation within a singe MLLM.
@@ 3.2 Unified Vision Tokenizer with Dual Codebooks
We utilize an Exponential Moving Average (EMA) strategy [36, 42] to update the codebook. To enhance codebook utilization, we implement a restart strategy where a cluster within the codebook is randomly replaced with an input from the current batch if it remains unused for a certain number of steps. To ensure the encoder outputs align closely with the codebook entries, we utilize a Vector Quantization (VQ) commitment loss, which is defined as: